# A Classification Theorem for Steady Euler Flows

### v1

@@ 4.1 The Unextendable Case
There exists $C>0$ such that for all $x\in\Omega_{\Gamma}$,
$\frac{dist(x,\partial\Omega_{\Gamma})^{2}}{C}\leq\psi(x)\leq Cdist(x,\partial\Omega_{\Gamma})^{2}.$ (4.6)
Now we are ready to apply the moving plane method to show that $\psi$ is even symmetric in the direction $e=(1,0)$. Without loss of generality, by translation, we fix $\inf_{x\in\Omega_{\Gamma}}x_{1}=0$. For $\lambda>0$, let $H_{\lambda}:=\{x_{1}=\lambda\}$. We now reflect $\psi$ and $\Omega_{\Gamma}$ through the hyperplane $H_{\lambda}$:
$\pi_{\lambda}(x_{1},x_{2}):=(2\lambda-x_{1},x_{2}),\ \psi_{\lambda}(x):=\psi(\pi_{\lambda}(x)),\ \Omega_{\lambda}:=\{x_{1}<\lambda\}\cap\Omega_{\Gamma},\text{ and }\breve{\Omega}_{\lambda}:=\pi_{\lambda}(\Omega_{\lambda}).$
@@
Since $\partial\Omega_{\Gamma}$ is analytic, for $\lambda\ll 1$, $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$. Moreover, as $\nabla\psi=0$ and $\partial_{nn}\psi=1$ in $\Omega_{\Gamma}$, $\psi$ satisfies the reflection monotonicity property at the level $0<\lambda\ll 1$ We will now prove that the reflection monotonicity property persists unless $\partial\breve{\Omega}_{\lambda}$ hits $\partial\Omega_{\Gamma}$ tangentially.
@@
$\lambda_{0}=\sup\{\lambda|\text{$\psi$ satisfies the reflection monotonicity property at level $\lambda$}\},$
@@
Figure 6: Internal tangency
Figure 7: Boundary tangency
The proof of Lemma 4.10 is based on the observation that as long as $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$, the difference $\psi-\psi_{\lambda}$ satisfies a singular elliptic equation where the singular coefficent has a uniform upper bound.
@@
For all $\lambda>0$, such that $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma},$ $h_{\lambda}:=\psi-\psi_{\lambda}$ satisfies a singular elliptic equation
$\Delta h_{\lambda}=c_{\lambda}(x)h_{\lambda},$ (4.7)
where the singular coefficient $c(x)$ has a uniform upper bound
$c_{\lambda}^{-}(x)\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}.$ (4.8)
@@
$\Delta h_{\lambda}=\frac{F(\psi)-F(\psi_{\lambda})}{\psi-\psi_{\lambda}}h_{\lambda}:=c_{\lambda}(x)h_{\lambda}.$ (4.9)
In the case where $F$ is analytic at $1$, as $F$ is already analytic in $(0,1)$, from Lemma 4.7, there exists a function $G\in C^{1}([0,1])$ and a constant $d$ such that $F(s)=d\sqrt{s}+G(s)$ ($d$ might be zero, which corresponds to the case where the Puiseux expression of $F$ at 00 has no $\sqrt{s}$ term). Now we have
$c_{\lambda}=\frac{d}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}$ (4.10)
Due to the lower bound of $\psi$ in the equation (4.6), we have
$\psi_{\lambda}>\frac{dist(x,\partial\breve{\Omega}_{\lambda})^{2}}{C},$ (4.11)
together with the fact that $\psi$ and $\psi_{\lambda}$ enjoy a uniform $C^{1}$ upper bound, we have $|c_{\lambda}|\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, where $M>0$ is independent of $\lambda$.
In the case where $F$ is not analytic near $1$, since $F$ is analytic in $(0,1)$, Lemma 4.5 and Lemma 4.7 show that there exists an integer $k_{0}\geq 2$, a negative constant $d_{1}<0$, a constant $d_{2}$ and a function $G\in C^{1}([0,1])$, such that $F(s)=d_{1}(1-s)^{\frac{k_{0}-1}{k_{0}}}+d_{2}\sqrt{s}+G(s)$. We now have the corresponding expression for $c_{\lambda}$
$c_{\lambda}=d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}+\frac{d_{2}}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}.$ (4.12)
As $d_{1}<0$, the term $d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}$ is non-negative. Hence, similiarly as in the case where $F$ is analytic at $1$, we have $c_{\lambda}^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, for a positive $M$ independent of $\lambda$. ∎
We now prove Lemma 4.10, arguing by contradiction.
@@
Due to the continuity of $\psi$, $h_{\lambda_{0}}:=\psi-\psi_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$. Now we assume that at the level of $\lambda_{0}$, none of the internal tangency or boundary tangency occurs, in particular, there is a $\delta^{*}>0$, such that for $\lambda_{0}\leq\lambda<\lambda_{0}+\delta^{*}$, $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$. We will show that there is a $\delta\in(0,\delta^{*})$ such that $\psi$ satisfies the reflection monotonicity at level $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$. We start by proving that $h_{\lambda_{0}}$ is positive in $\breve{\Omega}_{\lambda_{0}}$. By Lemma 4.11, $h_{\lambda_{0}}$ is a nonnegative solution to $\Delta h_{\lambda_{0}}=c(x)h_{\lambda_{0}}$ with $c^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}$. Now suppose that there is a point $x^{0}\in\breve{\Omega}$ such that $h_{\lambda}(x^{0})=0$, we automatically have $\nabla h_{\lambda}(x^{0})=0$. Then the Proposition A.1 in the appendix shows in a small ball whose boundary contains $x^{0}$, $h_{\lambda_{0}}$ is identically zero. Due to the analytic regularity of $\psi$, it further implies that $\psi_{\lambda_{0}}=\psi$. Hence, $\Omega_{\Gamma}$ is even symmetric to $H_{\lambda_{0}}$ and we get a contradiction. Hence, we now have
$h_{\lambda_{0}}(x)>0,\text{ for all $x\in\breve{\Omega}_{\lambda_{0}}$.}$ (4.13)
Anaglously, for any point $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$, we may apply Proposition A.1 to a small ball that hits $x$ boundary-tangentially at $x$, and conclude that
$\partial_{1}h_{\lambda_{0}}(x)>0,\text{ for all $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$},$ (4.14)
@@
In the rest of the proof, we will show that there is a $\delta>0$ such that $\psi$ satisfies the reflection monotonicity property at level $\lambda\in[\lambda_{0},\lambda+\delta].$ First, as $\breve{\Omega}_{\lambda_{0}}$ doesn’t hit $\partial\Omega_{\Gamma}$ boundary-tangentially, for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta^{*}]$, $H_{\lambda}\cap\partial\Omega_{\Gamma}$ consists of two points $x_{\lambda}^{1}$, $x_{\lambda}^{2}$($x_{\lambda}^{1}$ is above $x_{\lambda}^{2}$). Moreover, there is a $\epsilon_{0}>0$ such that $\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}}^{1})<-\epsilon_{0}$ and $\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}})(x_{\lambda_{0}}^{2})>\epsilon_{0}$. As $\partial_{12}h_{\lambda}(x_{\lambda}^{1})$ and $\partial_{12}h_{\lambda}(x_{\lambda}^{2})$ are uniformly continuous in $\lambda$ , we choose $\delta_{1}>0$ such that
$\partial_{12}h_{\lambda}(x_{\lambda}^{1})<-\frac{1}{2}\epsilon_{0},\partial_{12}h_{\lambda}(x_{\lambda}^{2})>\frac{1}{2}\epsilon_{0},\text{ for$\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta_{1}$}.$ (4.15)
@@
$\nabla h_{\lambda}(x_{\lambda}^{1})=0,$ (4.16)
moreover by symmetry, we have
$\partial_{11}h_{\lambda}(x_{\lambda}^{1})=\partial_{22}h_{\lambda}(x_{\lambda}^{1})=0.$ (4.17)
Due to the smoothness of $\partial\Omega_{\Gamma}$ and that the boundary tangency doesn’t occur when $\lambda=\lambda_{0}$, there exists a $\epsilon>0$, such that the direction $(1,-\epsilon)$ is always pointing to the exterior of $\breve{\Omega}_{\lambda}$ at $x_{\lambda}^{1}$ for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$(we may decrease the value of $\delta_{1}$ when it is necessary). Now since $D^{3}h_{\lambda}$ is uniformly bounded, from the equations (4.15), (4.16) and (4.17), there exists $r_{1}>0$, such that $h_{\lambda}>0$ in $B_{r_{1}}(x_{\lambda}^{1})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$. Analogously, in $B_{r_{1}}(x_{\lambda}^{2})$ (we decrease the value of $\delta_{1}$ when necessary), $h_{\lambda}\geq 0$ in $B_{r_{1}}(x_{\lambda}^{2})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$. Now due to the equation (4.14), there is a $\epsilon_{1}>0$ such that we have $\partial_{1}h_{\lambda_{0}}(x)>\epsilon_{1}$ for all $x\in H_{\lambda_{0}}\setminus(B_{r_{1}}(x_{\lambda_{0}}^{1})\cup B_{r_{1}}(x_{\lambda_{0}}^{2}))$. Similar to the previous case, we can find a $r_{2}>0$ and a $\delta_{2}\in(0,\delta_{1})$, such that $h_{\lambda}\geq 0$ in the set $T_{\lambda}:\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r_{2}\}\setminus(B_{r}(x_{\lambda}^{1})\cup B_{r}(x_{\lambda}^{2}))\}$ for all $\lambda_{0}\leq\lambda\leq\lambda+\delta_{2}$.
In summary, we show there is a uniform $r>0$ and $\delta>0$, such that if $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$, we have $h_{\lambda}\geq 0$ in $N_{\lambda}:=\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r\}$, which is a tubular neighborhood of $H_{\lambda}$. Note by (4.13) and Lemma 4.8, there is a $\epsilon_{2}>0$, such that $h_{\lambda_{0}}>\epsilon_{2}$ on the set $\breve{\Omega}_{\lambda_{0}}\setminus N_{\lambda_{0}}$. Now by the uniform continuity of $h_{\lambda}$, we may decrease the value of $\delta$ (still make it positive), such that $h_{\lambda}>\frac{\epsilon_{2}}{2}$ on the set $\breve{\Omega}_{\lambda}\setminus N_{\lambda}$ for all $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$. ∎

### v2

@@ 4.1. The Unextendable Case
There exists a constant $C>0$ such that for all $x\in\Omega_{\Gamma}$,
(4.5) $\frac{dist(x,\partial\Omega_{\Gamma})^{2}}{C}\leq\psi(x)\leq Cdist(x,\partial\Omega_{\Gamma})^{2}.$
Now we are ready to apply the moving plane method to show that $\psi$ is even symmetric in the direction $e=(1,0)$. Without loss of generality, by translation, we fix $\inf_{x\in\Omega_{\Gamma}}x_{1}=0$. For $\lambda>0$, define
$H_{\lambda}:=\{x_{1}=\lambda\},\,\,\Omega_{\lambda}:=\{x_{1}<\lambda\}\cap\Omega_{\Gamma},\text{ and }\pi_{\lambda}(x_{1},x_{2}):=(2\lambda-x_{1},x_{2}).$
We now reflect $\psi$ and $\Omega_{\lambda}$ through the hyperplane $H_{\lambda}$ as follows:
$\psi_{\lambda}(x):=\psi(\pi_{\lambda}(x)),\,\,\breve{\Omega}_{\lambda}:=\pi_{\lambda}(\Omega_{\lambda}).$
@@
Since $\partial\Omega_{\Gamma}$ is analytic, for $\lambda\ll 1$, $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$. Moreover, as $\nabla\psi=0$ and $\partial_{nn}\psi=1$ in $\partial\Omega_{\Gamma}$, $\psi$ satisfies the reflection monotonicity property at the level $0<\lambda\ll 1$ . We will now prove that the reflection monotonicity property persists unless $\partial\breve{\Omega}_{\lambda}$ hits $\partial\Omega_{\Gamma}$ tangentially.
@@
[…]
@@
Figure 4. Internal tangency and boundary tangency
The proof of Lemma 4.7 is based on the observation that as long as $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$, the difference $\psi-\psi_{\lambda}$ satisfies an elliptic equation where the coefficient has a uniform estimate.
@@
For all $\lambda>0$, such that $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma},$ $h_{\lambda}:=\psi-\psi_{\lambda}$ satisfies an elliptic equation
(4.6) $\Delta h_{\lambda}=c_{\lambda}(x)h_{\lambda},$
where the coefficient $c(x)$ satisfies the following uniform estimate
(4.7) $c_{\lambda}^{-}(x):=\max\{-c_{\lambda}(x),0\}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}.$
@@
(4.8) $\Delta h_{\lambda}=\frac{F(\psi)-F(\psi_{\lambda})}{\psi-\psi_{\lambda}}h_{\lambda}:=c_{\lambda}(x)h_{\lambda}.$
In the case where $F$ is analytic at $1$, as $F$ is already analytic in $(0,1)$, from Lemma 4.5, there exists a function $G\in C^{1}([0,1])$ and a constant $d$ such that $F(s)=d\sqrt{s}+G(s)$ ($d$ might be zero, which corresponds to the case where the Puiseux expansion of $F$ at 0 has no $\sqrt{s}$ term). Now we have
(4.9) $c_{\lambda}=\frac{d}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}$
It follows from the lower bound of $\psi$ in the estimate (4.5) that one has
(4.10) $\psi_{\lambda}\geq\frac{dist(x,\partial\breve{\Omega}_{\lambda})^{2}}{C}.$
This, together with the fact that $\psi$ and $\psi_{\lambda}$ enjoy a uniform $C^{1}$ upper bound, yields $|c_{\lambda}|\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, where $M>0$ is independent of $\lambda$.
In the case where $F$ is not analytic near $1$, since $F$ is analytic in $(0,1)$, Lemmas 4.4 and 4.5 show that there exist an integer $k_{0}\geq 2$, a negative constant $d_{1}<0$, a constant $d_{2}$, and a function $G\in C^{1}([0,1])$, such that
$F(s)=d_{1}(1-s)^{\frac{k_{0}-1}{k_{0}}}+d_{2}\sqrt{s}+G(s).$
We now have the corresponding expression for $c_{\lambda}$
(4.11) $c_{\lambda}=d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}+\frac{d_{2}}{\sqrt{\psi_{\lambda}}+\sqrt{\psi}}+\frac{G(\psi)-G(\psi_{\lambda})}{\psi-\psi_{\lambda}}.$
As $d_{1}<0$, the term $d_{1}\frac{(1-\psi)^{\frac{k_{0}-1}{k_{0}}}-(1-\psi_{\lambda})^{\frac{k_{0}-1}{k_{0}}}}{\psi-\psi_{\lambda}}$ is non-negative. Hence, similar to the case where $F$ is analytic at $1$, we have $c_{\lambda}^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda})}$, for a positive $M$ independent of $\lambda$. ∎
We now prove Lemma 4.7 by the contradiction argument.
@@
Due to the continuity of $\psi$, $h_{\lambda_{0}}:=\psi-\psi_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$. Now we assume that at the level of $\lambda_{0}$, none of the internal tangency or boundary tangency occurs, in particular, there is a $\delta^{*}>0$, such that for $\lambda_{0}\leq\lambda<\lambda_{0}+\delta^{*}$, $\breve{\Omega}_{\lambda}\subset\Omega_{\Gamma}$. We will show that there is a $\delta\in(0,\delta^{*})$ such that $\psi$ satisfies the reflection monotonicity at level $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$.
We start by proving that $h_{\lambda_{0}}$ is positive in $\breve{\Omega}_{\lambda_{0}}$. By Lemma 4.8, $h_{\lambda_{0}}$ is a nonnegative solution to
(4.12) $\Delta h_{\lambda_{0}}=c_{\lambda_{0}}(x)h_{\lambda_{0}}\,\,\text{with}\,\,c_{\lambda_{0}}^{-}\leq\frac{M}{dist(x,\partial\breve{\Omega}_{\lambda_{0}})}.$
Now suppose that there is a point $x^{0}\in\breve{\Omega}_{\lambda_{0}}$ such that $h_{\lambda_{0}}(x^{0})=0$, we automatically have $\nabla h_{\lambda_{0}}(x^{0})=0$. Then the Proposition A.1 in the appendix shows that in a small ball whose boundary contains $x^{0}$, $h_{\lambda_{0}}$ is identically zero. Due to the analytic regularity of $\psi$, it further implies that $\psi_{\lambda_{0}}=\psi$. Hence $\Omega_{\Gamma}$ is even symmetric with respect to $H_{\lambda_{0}}$ and we get a contradiction. Therefore, it holds that
(4.13) $h_{\lambda_{0}}(x)>0,\text{ for all $x\in\breve{\Omega}_{\lambda_{0}}$.}$
Analogously, for any point $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$, we may apply Proposition A.1 to a small ball that hits the boundary of $\partial\breve{\Omega}_{\lambda_{0}}$ tangentially at $x$, and conclude that
(4.14) $\partial_{1}h_{\lambda_{0}}(x)>0,\text{ for all $x\in H_{\lambda_{0}}\setminus\partial\Omega_{\Gamma}$},$
@@
In the rest of the proof, we will show that there is a $\delta>0$ such that $\psi$ satisfies the reflection monotonicity property at level $\lambda\in[\lambda_{0},\lambda_{0}+\delta].$ First, as $\breve{\Omega}_{\lambda_{0}}$ doesn’t hit the boundary $\partial\Omega_{\Gamma}$ tangentially, for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta^{*}]$, $H_{\lambda}\cap\partial\Omega_{\Gamma}$ consists of two points $x_{\lambda}^{1}$, $x_{\lambda}^{2}$($x_{\lambda}^{1}$ is above $x_{\lambda}^{2}$). Moreover, as $\breve{\Omega}_{\lambda_{0}}$ doesn’t hit the boundary $\partial\Omega_{\Gamma}$ tangentially at the points $x_{\lambda}^{1}$ and $x_{\lambda}^{2}$, the condition that $\partial_{nn}\psi=1,\psi_{tt}\psi=\psi_{tn}=0$ ($t,n$ here represent the tangent and normal direction, respectively) for all points in $\partial\Omega_{\Gamma}$ implies that there is an $\epsilon_{0}>0$ such that
$\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}}^{1})<-\epsilon_{0}\,\,\text{and}\,\,\partial_{12}h_{\lambda_{0}}(x_{\lambda_{0}})(x_{\lambda_{0}}^{2})>\epsilon_{0}.$
As $\partial_{12}h_{\lambda}(x_{\lambda}^{1})$ and $\partial_{12}h_{\lambda}(x_{\lambda}^{2})$ are uniformly continuous in $\lambda$, there exist a $\delta_{1}>0$ such that
(4.15) $\partial_{12}h_{\lambda}(x_{\lambda}^{1})<-\frac{1}{2}\epsilon_{0}\,\,\text{and}\,\,\partial_{12}h_{\lambda}(x_{\lambda}^{2})>\frac{1}{2}\epsilon_{0},\text{ for $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta_{1}$}.$
@@
(4.16) $\nabla h_{\lambda}(x_{\lambda}^{1})=0.$
Moreover, by symmetry, it holds that
(4.17) $\partial_{11}h_{\lambda}(x_{\lambda}^{1})=\partial_{22}h_{\lambda}(x_{\lambda}^{1})=0.$
Due to the smoothness of $\partial\Omega_{\Gamma}$ and that the boundary tangency doesn’t occur when $\lambda=\lambda_{0}$, there exists an $\epsilon>0$, such that the direction $(1,-\epsilon)$ is always pointing to the exterior of $\breve{\Omega}_{\lambda}$ at $x_{\lambda}^{1}$ for all $\lambda\in[\lambda_{0},\lambda_{0}+\delta]$(we may decrease the value of $\delta_{1}$ when it is necessary). Now since $\partial^{3}h_{\lambda}$ is uniformly bounded, from (4.15), (4.16), and (4.17), there exists an $r_{1}>0$ such that $h_{\lambda}>0$ in $B_{r_{1}}(x_{\lambda}^{1})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$. Analogously, $h_{\lambda}\geq 0$ in $B_{r_{1}}(x_{\lambda}^{2})$ for $\lambda\in[\lambda_{0},\lambda_{0}+\delta_{1}]$ (we decrease the value of $\delta_{1}$ when necessary). It follows from the estimate (4.14) that there is an $\epsilon_{1}>0$ such that we have $\partial_{1}h_{\lambda_{0}}(x)>\epsilon_{1}$ for all $x\in H_{\lambda_{0}}\setminus(B_{r_{1}}(x_{\lambda_{0}}^{1})\cup B_{r_{1}}(x_{\lambda_{0}}^{2}))$. Similar to the previous case, we can find $r_{2}>0$ and $\delta_{2}\in(0,\delta_{1})$, such that for all $\lambda\in(\lambda_{0},\lambda_{0}+\delta_{2})$, $h_{\lambda}\geq 0$ in the set
$T_{\lambda}:=\{x|x\in\breve{\Omega}_{\lambda},\,dist(x,H_{\lambda})<r_{2}\}\setminus(B_{r_{1}}(x_{\lambda}^{1})\cup B_{r_{1}}(x_{\lambda}^{2})).$
In summary, we show there are a uniform $r>0$ and $\delta>0$, such that if $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$, we have $h_{\lambda}\geq 0$ in $N_{\lambda}:=\{x|x\in\breve{\Omega}_{\lambda},dist(x,H_{\lambda})<r\}$, which is a tubular neighborhood of $H_{\lambda}$. Note by (4.13) and Lemma 4.6, there is an $\epsilon_{2}>0$, such that $h_{\lambda_{0}}>\epsilon_{2}$ on the set $\breve{\Omega}_{\lambda_{0}}\setminus N_{\lambda_{0}}$. Now by the uniform continuity of $h_{\lambda}$, we may decrease the value of $\delta$ (still make it positive), such that $h_{\lambda}>\frac{\epsilon_{2}}{2}$ on the set $\breve{\Omega}_{\lambda}\setminus N_{\lambda}$ for all $\lambda_{0}\leq\lambda\leq\lambda_{0}+\delta$. ∎
@@
Proof of Proposition 4.3.
Lemma 4.7 shows there exists a $\lambda_{0}>0$ with $h_{\lambda_{0}}\geq 0$ in $\breve{\Omega}_{\lambda_{0}}$ and at least one of the internal boundary tangency and boundary tangency happens for $\breve{\Omega}_{\lambda_{0}}$. In the case where internal tangency occurs at $x^{0}\in\partial\Omega_{\Gamma}$, by Lemma 4.8, $h_{\lambda_{0}}$ is a non-negative solution of the equation (4.12). Moreover, it holds that
$\nabla h_{\lambda_{0}}(x^{0})=\nabla\psi(x^{0})-\nabla\psi_{\lambda_{0}}(x^{0})=0.$
Now we apply Proposition A.1 for a sufficiently small ball $B\subset\breve{\Omega}_{\lambda_{0}}$ that hits $\breve{\Omega}_{\lambda_{0}}$ internal-tangentially at $x^{0}$, and conclude $h_{\lambda_{0}}=0$. Hence $\psi$ and correspondingly $\Omega_{\Gamma}$ is even symmetric in the direction $(1,0)$.
Now we consider the case where the boundary of tangency happens at $x^{0}\in H_{\lambda_{0}}\cap\partial\Omega_{\Gamma}$. First, similar to above, one has $\nabla h_{\lambda_{0}}(x^{0})=0$. In addition, $\partial\breve{\Omega}_{\lambda_{0}}$ forms a right angle at $x^{0}$, and the normal direction of $\partial\Omega_{\Gamma}$ at $x^{0}$ is parallel to direction $(0,1)$. In particular, we have $\nabla^{2}h_{\lambda_{0}}(x^{0})=0$. Now analogously as in the case of the internal tangency, by applying Proposition A.2 in the appendix to a small half ball contained in $\breve{\Omega}_{\lambda_{0}}$ that hits $\Omega_{\Gamma}$ boundary-tangentially, we have $h_{\lambda_{0}}=0$. Hence $\psi$ and correspondingly $\Omega_{\Gamma}$ is even symmetric in the direction $(1,0)$. ∎
4.2. The proof of Theorem 1.1
In this section, we finish the proof of Theorem 1.1.
We first introduce the following notation.
Definition 4.4.
Given a set $U\subset\mathbb{R}^{2}$ and a function $\psi\in C^{2}(U)$, if there exists a continuous function $F$ such that $\Delta\psi=F(\psi)$ in $U$, then $\psi$ is said to be realizable by $F$ in $U$, or we abbreviate as $\psi$ is realizable in $U$.
If $F$ is extendable in $\Omega_{\Gamma}$ as in Lemma 4.2, since $\Delta\psi-F(\psi)$ is an analytic function in a tubular neighborhood of $\partial\Omega_{\Gamma}$ and $\Delta\psi-F(\psi)=0$ in $\Omega_{\Gamma}$, $\psi$ must be realizable by $F$ in a domain containing $\Omega_{\Gamma}$. Furthermore, we have the following proposition.
Proposition 4.9.
Assume that $U$ is a simply connected bounded open set, $\psi\in C^{\omega}(\bar{U})$ is constant on $\partial U$}. Assume $\psi$ is realizable by $F$ in $U$ and $F$ is not analytic at $c:=\psi(\partial U)$, then $\psi$ is radially symmetric.